The resistance of a certain wire is 20 Ohms. Another wire of the same material has a diameter 1/3 as great and twice the length. What is the resistance of the second wire?
Answer #1

We know, that the resistance of a wire is given by

R=\rho\frac{L}{A},

where \rho is the resistivity of the material,

L is the length of the resistance, and

A is the area of its cross section.

It is given in the question that both wires are made of the same material. Therefore, we can conclude that the resistivities are equal for both of them.

We can thus say,

R_1=\rho\frac{L_1}{A_1}

and

R_2=\rho\frac{L_2}{A_2}

Dividing the two equations, we get

\frac{R_1}{R_2}=\frac{L_1}{L_2}\cdot \frac{A_2}{A_1}

Area of cross section is given by

A=\frac{\pi d^2}{4}

Thus, the equation derived above simplifies to

\frac{R_1}{R_2}=\frac{L_1}{L_2}\cdot \left(\frac{d_2}{d_1}\right) ^{2}

Putting L_2=2L_1 and d_2=\frac{d_1} {3} ​​​​​, we can solve for R_2 ​​​​​​.

\frac{20\:\Omega}{R_2}=\frac{L_1}{2L_1}\cdot \left(\frac{\frac{d_1}{3}}{d_1}\right) ^{2}

\Rightarrow R_2=360\:\Omega.​​​​​​